# RES-Q: Evaluating Code-Editing Large Language Model Systems at the Repository Scale

### v1

@@ Abstract
The instruction-following ability of Large Language Models (LLMs) has cultivated a class of LLM-based systems capable of approaching complex tasks such as making edits to large code repositories. Due to the high sensitivity and unpredictability of LLM behavior in response to changes in prompting, robust evaluation tools are needed to drive future iteration of these systems. We propose RES-Q, a natural language instruction-based benchmark for evaluating Repository Editing Systems, which consists of 100 repository editing tasks derived from real GitHub commits. Given an edit instruction and a code repository, RES-Q evaluates an LLM system’s ability to gather information and construct an edit that satisfies the criteria set by the instruction. We argue that evaluating LLMs in this way addresses issues with traditional benchmarks and provides a more holistic assessment of a model’s abilities. We evaluate various state-of-the-art LLMs as language agents in a repository-editing system built on Qurrent OS, our language agent development software. Despite their 1% pass@1 performance difference on HumanEval, we find Claude Sonnet 3.5 outperforms GPT-4o by 12% pass@1 on RES-Q, indicating RES-Q’s capacity to differentiate model capability as traditional benchmarks approach saturation. We further investigate token efficiency, performance relationships with existing benchmarks, and interesting disparities between closed and open-source LLMs. Code and dataset are available at this https URL.
@@ 1.3 Contributions
In this paper, we introduce RES-Q, a hand-crafted dataset consisting of 100 natural language code repository edit instructions paired with corresponding test suites. RES-Q’s instructions are derived from actual commits made to GitHub repositories, ensuring their relevance and realism. Concretely, RES-Q serves as a robust evaluation tool for assessing the performance of LLM-based systems on real-world software development tasks, particularly those involving the extension and modification of large code repositories. More generally, RES-Q provides an opportunity to evaluate the objective, binary success of an LLM working as a language agent on a complex task, thereby offering a more comprehensive indication of its capabilities compared to traditional benchmarks. To this end, we conduct an assessment of state-of-the-art models within a code-repository editing system built on Qurrent OS, our proprietary LLM agent development software, and compare the results with traditional benchmarks. To complement our dataset, we also release a submission environment to evaluate the correctness of repository changes for a given task, promoting broader community engagement.
@@
    Evaluation of SOTA LLMs: We evaluate state-of-the-art LLMs within a codebase-editing system built on Qurrent OS, our LLM agent development software, and investigate correlations with performance on traditional benchmarks (§ 4).
@@
    Submission Environment: We release a submission environment with a simple interface for evaluating the correctness of codebase changes for a given RES-Q task, facilitating the evaluation of other systems and encouraging broader community engagement.
@@ 4.2 Results
Constraining Context Window. For the selected closed-source models, constraining the context window size to 8K tokens consistently decreased performance on RES-Q. The largest performance decreases by constraining context were observed in the two top-performing models: Claude 3.5 Sonnet (-9%) and GPT-4o (-10%). For the selected open-source models, constraining the context window consistently increased performance. Qualitatively, this indicates that these models are not able to utilize their entire context window as effectively. Relative performance of models remained unchanged across the constrained and unconstrained settings.

### v2

@@ Abstract
The instruction-following ability of Large Language Models (LLMs) has cultivated a class of LLM-based systems capable of approaching complex tasks such as making edits to large code repositories. Due to the high sensitivity and unpredictability of LLM behavior in response to changes in prompting, robust evaluation tools are needed to drive future iteration of these systems. We propose RES-Q, a natural language instruction-based benchmark for evaluating Repository Editing Systems, which consists of 100 handcrafted repository editing tasks derived from real GitHub commits. Given an edit instruction and a code repository, RES-Q evaluates an LLM system’s ability to interpret the instruction, navigate the repository to gather relevant information, and construct an appropriate edit that satisfies the specified criteria. We argue that evaluating LLMs in this way addresses issues with traditional benchmarks and provides a more holistic assessment of a model’s abilities. We evaluate various state-of-the-art LLMs as language agents in a repository-editing system built on Qurrent OS, our language agent development software. Despite their 1% pass@1 performance difference on HumanEval, we find Claude Sonnet 3.5 outperforms GPT-4o by 12% pass@1 on RES-Q, indicating RES-Q’s capacity to differentiate model capability as traditional benchmarks approach saturation. We further investigate token efficiency, performance relationships with existing benchmarks, and interesting disparities between closed and open-source LLMs. Code and dataset are available at this https URL.
@@ 1.3 Contributions
In this paper, we introduce RES-Q, a hand-crafted dataset consisting of 100 natural language code repository edit instructions paired with corresponding test suites. RES-Q’s instructions are derived from actual commits made to GitHub repositories, ensuring their relevance and realism. Concretely, RES-Q serves as a robust evaluation tool for assessing the performance of LLM-based systems on real-world software development tasks, particularly those involving the extension and modification of large code repositories. More generally, RES-Q provides an opportunity to evaluate the objective, binary success of an LLM working as a language agent on a complex task, thereby offering a more comprehensive indication of its capabilities compared to traditional benchmarks. To this end, we conduct an assessment of state-of-the-art models within a code-repository editing system built on Qurrent OS, our proprietary language agent development software, and compare the results with traditional benchmarks. To complement our dataset, we also release a submission environment to evaluate the correctness of repository changes for a given task, promoting broader community engagement.
@@
    Evaluation of SOTA LLMs: We evaluate state-of-the-art LLMs within a codebase-editing system built on Qurrent OS, our language agent development software, and investigate correlations with performance on traditional benchmarks (§ 4).
@@
    Submission Environment: We release a submission environment with a simple interface for evaluating the correctness of codebase changes for a given RES-Q task, facilitating the evaluation of other systems and encouraging broader community engagement (§ 3.3).
@@ 4.2 Results
Constraining Context Window. For the selected closed-source models, constraining the context window size to 8K tokens consistently decreased performance on RES-Q. The largest performance decreases by constraining context were observed in the two top-performing models: Claude 3.5 Sonnet (-9%) and GPT-4o (-10%). For the selected open-source models, constraining the context window consistently increased performance. This suggests that these models may not be able to utilize their entire context window as effectively. Relative performance of models remained unchanged across the constrained and unconstrained settings.